OpenStudy (anonymous):

Which of these numbers is between 1 and 3/4? A.7/4 B.3/5 C.7/8 D.5/8

OpenStudy (moazzam07):

A

OpenStudy (moazzam07):

Because if 1 = 4/4 then 4/4 + 3/4 = 7/4

OpenStudy (apricot):

First, let's turn those fractions into decimals, to make it a lot easier. The answer has to be between .75 and 1. A.7/4 = 1.75, we can cross this out since it's over 1. B.3/5 = .6, we can also cross this out since it's under .75. C.7/8 = .875, this is your answer since it meets the requirements for the question. D.5/8 = .625, this is under .75.

OpenStudy (apricot):

C is your answer. :)